Firefox Fle Edit View Hstory Bookmarks Tools window Help 2i3. 0-???:) wed 9:49 AM David Frost a + E updates Raady to Install sonc-cocs could no, be updated 4726190&offset=next 200% Physical Chemistry Su 2018 David 2 Homework Set 4 Vapor Pressure and Polymer Solutions 2017-11-24 2 of 17 The molecular weight of a polymer is MP= 1.00x105 g/mol. Assume 100g of the polymer are dissoved in one literof water at T-298K. Calculate the mole fraction Xp of the polymer. Assume at T-298K water is 55.6M. ? Part A XP -1.79856-10-5 Previous Answers Correct
Answer #1

1) To calculate the molar fraction of the polymer in solution, the moles of polymer and water present in the system must be previously calculated.

Npol = Polymer Mass / MW Polymer = 100 gr / 1x10 ^ 5 gr / mol = 0.001 mol NH2O = Molarity H2O * Vol H2O = 55.6 mol / L * 1 L = 55.6 mol N Total = NPol + NH2O = 0.001 + 55.6 = 55.601 mol XPol = NPol / N Total = 0.001 / 55.601 = 1.7985 * 10 ^ -5 2) The respective molar volumes for each substance must be calculated: Vol Pol = 0.75 ml / gr * (1 * 10 ^ 5 gr / 1 mol) = 75000 mL / mol Vol H2O = 1 ml / gr * (18 gr / mol) = 18 ml / mol VMix: Sum of the molar volumes of each substance, multiplied by its respective molar composition: V mixture = (75000 * 1.7985 * 10 ^ -5) + (18 * (1-1.7985 * 10 ^ -5)) = 19.35 ml / mol 3) To calculate the volumetric fraction of the polymer, you must divide the volume of the polymer between the total volume: Op = (Vp / VT) = (75000 * 0.001 / 19.35 * 55.601) = 0.06971 4) Steam pressure solution follows the following equation: Pa = Paº * (1-Xb) where: Paº = Pressure of pure water vapor Pa = Steam pressure solution Xb = Mole fraction of solute Pa = 2535 Pa * (1-1.7985 * 10 ^ -5) = 2535.95 Pa.